# When the Correct Model Fails: The Optimality of Stackelberg Equilibria with Follower Intention Updates

### v2

@@ III-A Stackelberg dynamic game
Under FB information structure, the control chosen is $u^{i}_{t}~=~-K^{i}_{t}x_{t}$ for $i\in\{L,F\}$. We use $\Pi^{i}_{0:T-1}$ to denote player $i$’s strategy domain over the finite horizon [T1][T-1], and $\boldsymbol{\pi}^{i}\in\Pi^{i}_{0:T-1}$ to denote a specific strategy for player $i$. When the information structure context $\mathcal{I}^{i}$ is clear, we use $\boldsymbol{\pi}^{i}$ to denote $\boldsymbol{\pi}^{i}(\mathcal{I}^{i})$.
@@ IV-A Open loop Stackelberg equilibrium (OLSE) under BR updates
Proposition 1 provides a sufficient condition for when the true follower’s BR belief $b^{\star}$ minimizes the remaining cost. The condition $x^{1}_{\tau}=x^{\star}_{\tau}$ is quite restrictive in general, and is false for a belief $b^{1}\neq b^{\star}$, $b^{1}\in\mathcal{F}$. When $x^{1}_{\tau}\neq x^{\star}_{\tau}$, then it is possible that $J^{L}_{\tau:T}(b^{2})<J^{L}_{\tau:T}(b^{\star})$.
@@
The leader first calculates the strategies $\mathbf{u}^{L\star}$ and $\mathbf{u}^{L^{\prime}}$ under beliefs $b^{\star}$ and $b^{\prime}$ (numerical values in App. -B). At update time $\tau$, the leader’s cost under both BR beliefs is
@@
At $t=\tau$ the state under BR beliefs $b^{\star}$ is $x^{\star}_{3}~=~1.22$ and the state under BR belief $b^{\prime}$ is $x^{\prime}_{3}~=~2.13$. The updated leader controls are $\hat{u}^{L\star}_{3:4}$ and $\hat{u}^{L^{\prime}}_{3:4}$ (App. -B). We verify that the OLSE of this problem is time-inconsistent since $u^{L\star}_{3:4}~\neq~\hat{u}^{L\star}_{3:4}$.
Finally, the cost for the horizon $[3,5]$ under belief $b^{\star}$ and $b^{\prime}$ is
@@
Despite the cost for the horizon $[3,5]$ being greater for $b^{\prime}$, this belief still obtains lower total cost over $[T]$ since $J^{L}_{0:5}(b^{\star})~=~1493.9$ and $J^{L}_{0:5}(b^{\prime})~=~1390.6$.
@@ IV-B Feedback Stackelberg equilibrium (FSE) under BR updates
with a horizon $T=8$, update time $\tau=3$, and initial condition $x_{0}=-5.6$. The leader’s cost (42) is defined by the matrices $(Q^{L},R^{L})=(7,16)$. The follower’s true BR $b^{\star}$ is given by $(Q^{F},R^{F})=(4,24)$. We consider the case where the leader also assumes a BR $b^{1}=b^{2}=b^{\prime}$ defined by $(Q^{F},R^{F})=(29,12)$. Because the FSE is time-consistent and the beliefs remain constant, the sequence of leader gains under beliefs $b^{\star}$ and $b^{\prime}$, respectively, are $\mathbf{K}^{L\star}$ and $\mathbf{K}^{L^{\prime}}$ and the follower’s BR in either case is $\mathbf{K}^{F\star}$ and $\mathbf{K}^{F^{\prime}}$. All numerical values are provided in App. -C. After playing the different strategies, the leader observes costs
@@ V-D Cost difference
% $\%\text{ higher cost}=\frac{J^{L}_{0:T}(\hat{b})-\min_{b}J^{L}_{0:T}(b)}{\min_{b}J^{L}_{0:T}(b)}\cdot 100$ (41)

### v3

@@ III-A Stackelberg dynamic game
Under FB information structure, the control chosen is $u^{i}_{t}~=~-K^{i}_{t}x_{t}$ for $i\in\{L,F\}$. We use $\Pi^{i}_{0:T-1}$ to denote player $i$’s strategy domain over the finite horizon [T1][T-1], and $\boldsymbol{\pi}^{i}\in\Pi^{i}_{0:T-1}$ to denote a specific strategy for player ii. When the information structure context $\mathcal{I}^{i}$ is clear, we use $\boldsymbol{\pi}^{i}$ to denote $\boldsymbol{\pi}^{i}(\mathcal{I}^{i})$.
@@ IV-A Open loop Stackelberg equilibrium (OLSE) under BR updates
Proposition 1 provides a sufficient condition for when the true follower’s BR belief $b^{\star}$ minimizes the remaining cost. The condition $x^{1}_{\tau}=x^{\star}_{\tau}$ is quite restrictive, and in general may be false for a belief $b^{1}\neq b^{\star}$, $b^{1}\in\mathcal{F}$. When $x^{1}_{\tau}\neq x^{\star}_{\tau}$, then it is possible that $J^{L}_{\tau:T}(b^{1})<J^{L}_{\tau:T}(b^{\star})$.
@@
The leader first calculates the strategies $\mathbf{u}^{L\star}$ and $\mathbf{u}^{L^{\prime}}$ under beliefs $b^{\star}$ and $b^{\prime}$, given by
$\mathbf{u}^{L\star}$ $=\left[-4.6,-0.86,0.06,0.19,0.12\right],$ (23)
$\mathbf{u}^{L^{\prime}}$ $=\left[-2.28,0.4,0.73,0.52,0.25\right].$ (24)
At update time $\tau$, the leader’s cost under both BR beliefs is
@@
At $t=\tau$ the state under BR beliefs $b^{\star}$ is $x^{\star}_{3}~=~1.22$ and the state under BR belief $b^{\prime}$ is $x^{\prime}_{3}~=~2.13$. The updated leader controls are $\hat{u}^{L\star}_{3:4}$ and $\hat{u}^{L^{\prime}}_{3:4}$ given by
$\hat{u}^{L\star}_{3:4}=\left[-1.06,-0.3\right],\quad\hat{u}^{L^{\prime}}_{3:4}=\left[-1.56,-0.41\right].$ (26)
We verify that the OLSE of this problem is time-inconsistent since $u^{L\star}_{3:4}~\neq~\hat{u}^{L\star}_{3:4}$.
Finally, horizon $[3,5]$ incurs cost under belief $b^{\star}$ and $b^{\prime}$ as
@@
Despite the cost for the horizon $[3,5]$ being greater for $b^{\prime}$, this belief still obtains lower total cost over $[T]$ since $J^{L}_{0:5}(b^{\star})~=~1493.9$ and $J^{L}_{0:5}(b^{\prime})~=~1390.6$. Additional numerical values for this example are given in App. -B.
@@ IV-B Feedback Stackelberg equilibrium (FSE) under BR updates
with a horizon $T=8$, update time $\tau=3$, and initial condition $x_{0}=-5.6$. The leader’s cost (49) is defined by the matrices $(Q^{L},R^{L})=(7,16)$. The follower’s true BR $b^{\star}$ is given by $(Q^{F},R^{F})=(4,24)$. We consider the case where the leader also assumes a BR $b^{1}=b^{2}=b^{\prime}$ defined by $(Q^{F},R^{F})=(29,12)$. Because the FSE is time-consistent and the beliefs remain constant, the leader strategy under beliefs $b^{\star}$ and $b^{\prime}$, respectively, are $\mathbf{K}^{L\star}$ and $\mathbf{K}^{L^{\prime}}$ and the follower’s BR in either case is $\mathbf{K}^{F\star}$ and $\mathbf{K}^{F^{\prime}}$, all control numerical values are given by
$\mathbf{K}^{L\star}=$ $[0.31,0.31,0.31,0.31,0.32,0.32,0.33,0.3],$ (34)
$\mathbf{K}^{L^{\prime}}=$ $[0.01,0.01,0.01,0.01,0.01,0.01,0.01,0.02],$ (35)
$\mathbf{K}^{F\star}=$ $[0.27,0.27,0.27,0.26,0.26,0.26,0.25,0.19],$ (36)
$\mathbf{K}^{F^{\prime}}=$ $[0.37,0.37,0.37,0.37,0.36,0.35,0.31,0.19].$ (37)
After playing the different strategies, the leader observes costs
@@ V-D Cost difference
% $\%\text{ higher cost}=\frac{J^{L}_{0:T}(\hat{b})-\min_{b}J^{L}_{0:T}(b)}{\min_{b}J^{L}_{0:T}(b)}\times 100.$ (48)